the velocity of the particle noving in the x-y plane is given by dx/dt=8asin2at,dy/dt=5acos2at.when t=0,x=8andy=0,the path of the particle is a.................here a=3.14
sir plz tell some 2 way to find path of particle

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear jauneet

to find the path means you have to find the ralation ship between x and y with the given information .

here we have to eleminate the t .

dx/dt=8asin2at

intigrate

 x= -(8acos2at )/2a   + c1

 x = -4cos2at + c1

now use the given condition

at t=0 ,x=8

so 8=-4 +c1

 c1=12

 so equation become

x =-4cos2at + 12 

or  (x-12)/4 = -cos2at ............1

now similerly for y

dy/dt=5acos2at

intigrate

y= (5asin2at)/2a  +c2

y =5/2 sin2at + c2

now apply condition

y=0 at t=0

0 =0+c2

c2=0

so equation become

y =5/2 sin2at

or 2y/5 = sin2at ............2

 

now square equation 1 and 2 and add

((x-12)/4)2 + (2y/5)2 = 1

this the path of the article

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed solution very  quickly.

 We are all IITians and here to help you in your IIT JEE  preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin